OpenStudy (anonymous):

The point (12, 9) is included in a direct variation. What is the constant of variation?

OpenStudy (anonymous):

the formula for direct variation is y=kx. The k is the constant. If you put the x and y values from the ordered pair and solve for k, you'll have the constant. 9=12k Divide both sides by 12 9/12=k simplify k=3/4

OpenStudy (anonymous):

direct variation = y = kx k is the constant variation Plugin the points 9/12 = 12k/12 3/4 = k k = 3/4